The time it takes to hand carve a guitar neck is uniformly distributed between 104 and 152 minutes. Determine the standard deviation of the completion time for carving the guitar neck.
Answer #1

Solution :

Given that,

a = 104

b = 152

standard deviation =  (b - a)2 / 12 = (152- 104)2 / 12 = 13.856